OpenStudy (anonymous):

hey um what's the radus of the base of a cylinder that has a height of 8cm and a volume of 628cm3 and round your answer to 0.1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

r=5

OpenStudy (anonymous):

Okay, the volume of a cylinder is given by \[\pi r^2 h\] so set \[628=\pi r^2 (8) \] Then solve for r.

OpenStudy (anonymous):

how do u solve that